In order to determine the emf of a storage battery it was conneted in series with a standard cell in a certain circuit and a current I1was obtained. When the battery is connected to the same circuit opposite to the standard cell a current I2 flows in 
the external circuit leaving from the positive plate of the storage battery. What is the emf of the standard cell is 
of the storage battery ? The emf of the standard cell is E2.
 
 
 

Grade:12

6 Answers

Sher Mohammad IIT Delhi
9 years ago
E=standard+Emf = V+ir
-v-ir+Estandard= Emf
v2+Ir= E
solving above 2 give the standard emf.
Sher Mohammad IIT Delhi
9 years ago
let asuume E is the unknown potential
E+E2= V1+I1R
E= v2+I2R
 
taking differnce 
E2= v1-v2+ (I1-I2)R
This equation give R, 
putting back R gives E , the unkwon emf.
 
sher mohammad
faculty askiitians 
b.tech, iit delhi
 
Sher Mohammad IIT Delhi
9 years ago
let asuume E is the unknown potential
E+E2= V1+I1R
E= v2+I2R
 
taking differnce 
E2= v1-v2+ (I1-I2)R
This equation give R, 
putting back R gives E , the unkwon emf.
 
sher mohammad
faculty askiitians 
b.tech, iit delhi
 
Sher Mohammad IIT Delhi
9 years ago
let asuume E is the unknown potential
E+E2= V1+I1R
E= v2+I2R
 
taking differnce 
E2= v1-v2+ (I1-I2)R
This equation give R, 
putting back R gives E , the unkwon emf.
 
sher mohammad
faculty askiitians 
b.tech, iit delhi
 
Sher Mohammad IIT Delhi
9 years ago
let asuume E is the unknown potential
E+E2= V1+I1R
E= v2+I2R
 
taking differnce 
E2= v1-v2+ (I1-I2)R
This equation give R, 
putting back R gives E , the unkwon emf.
 
sher mohammad
faculty askiitians 
b.tech, iit delhi
 
Kaushal
4 years ago
 
let asuume E is the unknown potential
E+E2= V1+I1R
E= v2+I2R
 
taking differnce 
E2= v1-v2+ (I1-I2)R
This equation give R, 
putting back R gives E , the unkwon